A vitamin syrup for infants contains 27.44 mg of zinc sulfate monohydrate (ZnSO4.H2O) and 100.00 mg of ascorbic acid (C6H8O6)for every 5 ml.

How many molecules of zinc sulfate monohydrate does a 5 ml of vitamin have?

Note: Round off your answer to two decimal places.
1
Expert's answer
2020-10-22T14:18:18-0400

n = m/M

m (ZnSO4.H2O) = 27.44/1000 = 0.02744 g

M (ZnSO4.H2O) = 179.49 g/mol

n (ZnSO.H2O) = 0.02744/179.49 = 0.00015 mol

Each mole of matter contains 6.02214076×1023 molecules.

Given amount of ZnSO.H2O is equal to: 0.00015 x 6.02214076×1023 = 1020 molecules.